# Truthful Text Sanitization Guided by Inference Attacks

### v1

@@ 5.3 Baselines
    Presidio’s synthetic replacement [^9] https://github.com/microsoft/presidio/blob/main/docs/samples/python/synth_data_with_openai.ipynb: replaces the text span entity type with a synthetic alternative (e.g., “[LOCATION]$\rightarrow$“Spain”) generated with GPT-3.5-turbo As mentioned in Section 2.2, because replacements are not intended to be truthful, they are expected to notoriously distort the document’s meaning.
@@
    Dou et al. replacement: replaces the text span by one of the three alternative replacements suggested by the model from [13]. [^10] For each text span, one of the three candidates is randomly chosen (as the three suggestions of this model are not ordered). The exact model can be accessed at https://huggingface.co/douy/Llama-2-7B-lora-instruction-ft-abstraction-three-span.
@@ 5.4 General statistics
Figure 2 shows the frequency of which candidate was finally selected with the INTACT approach from the list of five replacement candidates on the TAB corpus. The candidates are sorted by the LLM by order of abstraction, from the most specific to the most generic/abstract. In a little over half of the cases (53%), the most specific (first) replacement candidate was chosen [^11] For DATETIME entities, the frequent selection of Candidate 1 and 2 can be attributed, in part, to the stricter definition of what constitutes a ”successful” guess, where only complete lemma overlap was considered as guessed..
In about 22% of text spans in the TAB corpus, all five replacement candidates were guessed by the LLM and the entity label was thus used to replace the text span. This default replacement strategy was particularly common for entities labeled as ORG and MISC, amounting to almost half of text spans ending up in this default case, as detailed in Table 5.
@@ 5.5 Utility evaluation
For evaluating utility preservation, we first apply the TPS metric proposed in Section 4.1 between the original and protected documents. [^12] To determine the value for the hyper-parameter $N$ specifying the spacing between text spans replaced by [MASK] tokens (see Section 4.1.1), we conducted tests across multiple values (i.e., $N\in[2,4,6,8]$), and took $N=6$ as a good trade-off between runtime and the considered context size. Figure 3 depicts the TPS values obtained for the different strategies. Since all of them protected the same text spans (that is, those manually annotated in TAB), discrepancies in results arise from the varying semantic similarities between the original text spans and their replacements.
Figure 3: TPS for the different replacement strategies
The worst utility corresponds to the suppression strategy, whereas utility figures strictly increase as we employ more specific/utility-preserving replacements, that is: entity types $\rightarrow$ Presidio replacements $\rightarrow$ least-specific replacements $\rightarrow$ Dou et al. replacements $\rightarrow$ INTACT replacements $\rightarrow$ most-specific replacements. As a matter of fact, the average semantic similarities for the different replacements were 0.33, 0.38, 0.49, 0.49, 0.51, and 0.62, respectively.
As mentioned in Section 5.1, about 11% of the text spans in TAB were marked by the annotators as expressing direct or indirect PII. The suppression of those PII-related text spans causes a 22% drop in TPS, suggesting these spans are, on average, twice as informative as the non-protected spans. As illustrated in Figure 2, the two most common replacement candidates selected by INTACT are respectively the first candidate of the sorted list of replacements or the default entity type, which explains why its TPS falls between that of entity replacements (i.e., 85) and the most-specific replacements (i.e., 92).
@@
To further validate the results, we also evaluated the utility preserved by the different strategies on a downstream task: document clustering. To quantify the task-oriented utility preservation, we compared the clusters obtained from the original documents (which we considered the ground truth), with those resulting from the protected documents under the different strategies. Utility preservation was then computed as the similarity between both clusterings, as similar clusters suggest that equally similar conclusions can be reached from the analysis of the protected documents.
@@
To ensure that clustering leverages the semantic content of the document, we followed the method from [42], which relies on BERT [^13] More specifically the bert-base-cased model, see https://huggingface.co/google-bert/bert-base-cased. [12] to generate document embeddings by averaging the embeddings over all individual tokens. Once the document embeddings were obtained, they were clustered using the K-means++ algorithm [6] with $K=4$. The value of $K$ was selected based on the optimal inertia achieved when clustering the original documents. To mitigate the effect of the random initialization, the clustering for each document set was repeated 50 times, choosing the results with the lowest inertia. To quantify the similarity between clustering results, we used the Normalized Mutual Information (NMI) [14]. For NMI calculation, the average of 5 runs of the aforementioned clustering process was computed. In summary, each clustering was performed 250 times ($50\times 5$) to minimize the impact of the random initialization.
@@
Clustering-based utility evaluation for the considered strategies are shown in Figure 4. If we compare the results of this clustering-based utility evaluation with TPS, we obtain a strong Pearson correlation $r=0.864$ ($p$-$value=0.012$). This shows the effectiveness of TPS as an a priori estimator of the analytical utility of the protected documents in downstream tasks.
Figure 4: NMI for the clustering resulting from the evaluated replacement strategies
@@
Method-level differences offer additional insights into how various approaches impact data utility. For instance, as one can observe from Figure 4, entity replacements obtained a higher NMI than least-specific replacements, while the opposite holds for TPS. Similarly, INTACT achieves a NMI higher to that of the most-specific replacements, yet their TPS is lower. A common factor in both cases is the use of entity types. Although generated replacements are more similar to the original text spans when examined individually (i.e., higher TPS), some tend to distort the general document content more than entity types (i.e., lower NMI). Such distortion likely arises when a generalization leads to overly broad replacements that disrupt the linguistic register and structure of the document (as exemplified by the three least-specific replacements in Example 1), or even unfaithfully abstract its initial content. The strategy of using entity types when all generated replacements are guessed (which is frequent, as shown in Figure 2) in fact resulted in a NMI that surpassed that of the most-specific replacements. It seems that, when the language model fails to generate replacements that sufficiently protect privacy, it also struggles to produce options that effectively preserve utility.
@@
Discrepancy between TPS and NMI is significantly more pronounced with Dou et al. and least-specific replacements, which achieved a TPS nearly identical to INTACT but resulted in a much lower NMI. The replacements for these methods are observed to often be lengthy sentences that, while resembling the original span (i.e., high TPS), tend to disrupt more the overall text meaning (i.e., low NMI).
@@
Presidio’s synthetic replacements achieved a notably high NMI, outperforming both entity-based and least-specific replacements, likely due to their better integration within the text. Nonetheless, even when using a more advanced model than the one used to generate our replacements (i.e., GPT-3.5-turbo $>$ Mistral-7B [^14] https://lmarena.ai/?leaderboard), its non-truthful replacements still yielded a lower NMI compared to our most-specific and INTACT’s selected replacements.
@@ 5.6 Privacy evaluation
We evaluated the residual disclosure risk of the strategies by using the TRIR metric introduced in Section 4.2161616For TRIR calculation, we leveraged the ready-to-use implementation available in the corresponding GitHub repository  [^15] https://github.com/BenetManzanaresSalor/TextRe-Identification with the default attack hyperparameters..
The use of TRIR requires the specification of a body of background knowledge on the basis of which the re-identification risk is derived. We relied on a corpus of 13,759 rulings published on the ECHR website to define this background knowledge. The annotated court decisions from TAB were derived from the same source, this corpus constitutes therefore a strict superset of the TAB documents. However, assuming that each adversary has access to this full corpus of rulings is somewhat unrealistic, as such an adversary would already have access to all original documents, and would therefore have no need to re-identify the sanitized versions of those same documents. On the other hand, considering that the knowledge of the adversary strictly encompasses the protected individuals is also unrealistic, as they would rarely known the exact population sample in the protected dataset. Taking this into consideration, we defined the background set as a random superset of TAB encompassing twice the number of individuals/documents (i.e., $1,268\times 2=2,536$), and considering the first 10% of non-Procedure text from the rulings.
@@
Figure 5: TRIR for different replacement strategies
As expected, the replacements with the most-specific LLM generalisations yield the worst TRIR scores. The difference between this approach and the suppression strategy is, however, relatively small (2.2 percent points). In other words, the reliance on informative replacements (generalisations) makes it possible to substantially improve utility without substantially impairing privacy with respect to the strictest protection, i.e., suppression. INTACT performed particularly well on this balance, achieving a relatively low TRIR overall – second lowest when considering only generated replacements –, while delivering the second-highest TPS and the highest NMI. Compared to both the least-specific and most-specific replacements, INTACT delivered lower TRIR alongside comparable or superior TPS and NMI.
On the other hand, Presidio’s and entity replacements incurred the lowest risks because of the non-truthful replacements of the former (see Section 5.7), and the abstract replacements of the latter tended to confuse the re-identification model. This reduced the re-identification accuracy even more than suppression, because the re-identification model relies on concepts found in the background knowledge of the individuals targeted for re-identification. While suppression removes part of that exploitable information, Presidio’s synthetic and entity replacements substitute it by new information that is rarely found in any background knowledge. The scenario differs for truthful replacements –like INTACT or those from Dou et al.–, where part of the meaning of the original concept is preserved, and therefore it is possible to connect it with the corresponding individual’s background knowledge, which sightly increases TRIR.
However, regardless of the actual TRIR, using non-truthful replacements as Presidio does, may allow an attacker to differentiate between protected and non-protected terms, thereby providing an advantage that can be exploited to conduct case-specific attacks focused on inferring only the protected terms. Also, an obviously protected text, as it is the case for entity replacements, which use abstract terms, raises awareness of the document’s sensitivity and attracts more attention as a target for attacks [8].
@@ 5.7 Truthfulness and abstraction level
An important characteristic of the INTACT approach is its emphasis on producing truth-preserving replacements whose semantic content subsumes the content of the original text span. Such replacements should therefore correspond to generalisations that encompass and abstract away the original text span. However, the truthfulness and abstraction level of the replacements are nuances that are challenging to evaluate automatically in a reliable manner. Their impact in combination is reflected in the results of our automatic metrics, but these do not provide explicit information about replacement truthfulness and abstraction level separately.
@@
To address this, we conducted a manual evaluation focused on the truthfulness and abstraction level of the selected replacements. This evaluation was performed on a subset of TAB documents, and comprised 600 instances, each consisting of a PII span mention with 70 characters preceding and following it as context. For each document, we randomly sample a mention for each entity and then select a random subset of all these sampled mentions, using an equal size of 100 instances per each entity type, excluding the PERSON and CODE categories. The instances were rated by a single evaluator with NLP background, who was also allowed to look up concepts on the internet.
@@
The selected replacements are assessed using two criteria: abstraction level and truthfulness. A replacement is considered more abstract if it constitutes a hypernym or provides fewer details than the original PII span. A replacement is truthful if it retains the same overall content of the original span. When approximate equivalence can be assumed in the context of the span (e.g.28 days and 1 month), it is counted as truth-preserving. Partially true spans (e.g., containing two dates) are not considered truthful.
@@
Table 1 presents the manual evaluation results for Presidio, Dou et al.’s model [13] and INTACT. According to this manual evaluation, 81.33% of the assessed replacements were less specific and 93.17% truthful with INTACT. In contrast, Presidio’s manual evaluation reports 36.16% less specific and just 19.67% truthful replacements. Dou et al.’s abstraction model also produced mostly more abstract and mostly truthful replacements (78.33% and 95.16% respectively). However, replacements that were truthful but not more abstract amounted to 20.16% for Dou et al., which is 5% higher than for INTACT, where the iterative inference attacks are meant to prevent the use of too specific replacements. This supports the TRIR-based automatic evaluation results (Figure 5) which show a higher privacy risk posed by Dou et al. than INTACT.
@@
Table 1: Manual evaluation of replacements on truthfulness and abstraction level for Presidio, Dou et al and INTACT.
Approach Not-truthful Truthful
@@
Presidio 61.83% 18.5% 2% 17.67%
Dou et al. 1.5 % 3.33% 20.16% 75%
INTACT 3.5% 3.33% 15.17% 78%
The results from Table 1 show that smaller, open LLMs are well suited to finding generalized replacements in terms of abstraction level and, especially, truthfulness, even without specific fine-tuning for this task. For both INTACT and Dou et al., however, the generation of more abstract replacements was somewhat more challenging to achieve for the DEM and the MISC categories, where text spans are more often specialized or technical terms. The LLMs showed some tendency to produce definitions, explanations or paraphrases for these terms. Sometimes there was still a degree of abstraction between the different alternatives, but in some cases, none of the suggestions were generalizations, e.g., unemployed resulted in the INTACT replacements: Out of work, Jobless, Inactive in the labor force, Economically inactive, Without employment. We also noticed that the truthful and more abstract replacements obtained with Dou et al. were often more generic (e.g., in a specific period for replacing a DATETIME entity) than the ones for INTACT. This type of semantic loss is likely one of the reasons why INTACT outperformed Dou et al. in the clustering-based utility evaluation results presented in Figure 4.
In comparison, Presidio’s replacements were much less consistent, with a vast majority of them being non-truthful and with arbitrary degrees of abstraction. The lack of truthfulness and consistency is expected to significantly distort the document’s meaning. Moreover, we noticed a substitution bias towards the American-related sources, with the nationality of defendants in the TAB corpus changed from Turkish, Polish and British to Brazilian, Canadian and Mexican. Even though these distortions contribute to preventing re-identification (as shown in Section 5.6), it also seriously hampers secondary uses, as it might invalidate conclusions drawn from per-document analysis.

### v2

@@ 5.3 Baselines
[…]
@@
    Dou et al. replacement [18]: replaces the text span by one of the three alternative replacements provided by a fine-tuned model. [^11] For each text span, one of the three candidates is randomly chosen (as the three suggestions of this model are not ordered). The exact model can be accessed at https://huggingface.co/douy/Llama-2-7B-lora-instruction-ft-abstraction-three-span.
@@ 5.4 General statistics
Figure 4 shows the frequency at which each of the five replacement candidates is selected by the INTACT approach. The candidates are sorted by the LLM by order of abstraction, from the most specific to the most generic/abstract. In 53% of the text spans from the TAB corpus, the most specific (first) replacement candidate is chosen [^12] For DATETIME entities, the frequent selection of Candidate 1 and 2 can be attributed, in part, to the stricter definition of what constitutes a “successful” guess, where only complete lemma overlap was considered as guessed.. In about 22% of the text spans , all five replacement candidates were guessed by the LLM and the entity label was thus used to replace the text span. This default replacement strategy was particularly common for entities labeled as ORG and MISC, amounting to almost half of text spans ending up in this default case, as detailed in Table 5.
The experiments were run locally on a single NVIDIA RTX 4000 SFF Ada GPU with 20GB of memory. The candidate generation step took approximately 3 hours for the 127 documents containing over 7,300 entity mentions in the TAB test set, while inference attacks took considerably longer (approximately 15-20 times more) with the smaller, quantized open LLMs tested. It is worth noting that LLM calls for replacement selection are performed for each replacement candidate for each entity mention based on a whole document. In contrast, generation is performed once per each unique entity.
@@ 5.5 Utility evaluation
Utility preservation is examined from both intrinsic and extrinsic viewpoints, as detailed below.
@@
We first assess the preservation of the document’s semantic content by comparing the original documents and their protected versions using the application-independent TPS metric proposed in Section 4.1131313To determine the value for the hyper-parameter $N$ specifying the spacing between text spans replaced by [MASK] tokens (see Section 4.1.1), we tested multiple values (i.e., $N\in[2,4,6,8]$), and selected $N=6$ as a trade-off between runtime and context size. Figure 5 depicts the TPS values obtained for the different strategies. Since all of them protected the same text spans (manually annotated in TAB), discrepancies in TPS scores reflect the varying semantic similarities between the original text spans and their replacements.
Figure 5: Text Preserved Similarity (TPS) for the different replacement strategies (higher is better). The three last results, shown in green, all leverage the Mistral 7B instruct model.
The lowest utility corresponds to the suppression strategy, whereas utility figures strictly increase as we employ more specific/utility-preserving replacements, that is: entity types $\rightarrow$ Presidio replacements $\rightarrow$ least-specific replacements $\rightarrow$ Dou et al. replacements $\rightarrow$ INTACT replacements $\rightarrow$ most-specific replacements. The average semantic similarities for the different replacements were 0.33, 0.38, 0.49, 0.49, 0.51, and 0.62, respectively.
As mentioned in Section 5.1, about 11% of the text spans in TAB were marked by the annotators as expressing direct or indirect PII. The suppression of those PII-related text spans causes a 22% drop in TPS, suggesting these spans are, on average, twice as informative as the non-protected spans. As shown in Figure 4, the two most common replacement candidates selected by INTACT are respectively the first candidate of the sorted list of replacements or the default entity type, which explains why its TPS falls between that of entity replacements (85%) and the most-specific replacements (92%). In general, replacement-based methods show TPS differences less than 5%. This is likely due to the fact that they share a common set of text spans as replacement candidates, with the broadest generalization (entity type) already resulting in TPS=85%.
@@
To further validate the results, we measured to which extent document’s utility was preserved by the different strategies on a downstream task: document clustering. To quantify the task-oriented utility preservation, we compared the clusters obtained from the original documents (viewed as the ground truth), with those resulting from the protected documents under the different strategies. Utility preservation was then computed as the similarity between both clusterings, as similar clusters suggest that equally similar conclusions can be reached from the analysis of the protected documents. Document clustering, being a broad and unsupervised task based on the similarities of documents, serves as a good proxy for how the sanitized texts could perform in a variety of applications. It is particularly useful because it does not rely on predefined labels, thereby reflecting how well semantic and topical structures are preserved post-sanitization.
@@
The clustering method followed the approach proposed in [60], which relies on BERT [16] [^14] More specifically the bert-base-cased model, see https://huggingface.co/google-bert/bert-base-cased. to generate document embeddings by averaging the embeddings over all individual tokens. Once the document embeddings were obtained, they were clustered using the K-means++ algorithm [7] with $K=4$. The value of $K$ was selected based on the optimal inertia achieved when clustering the original documents. To mitigate the effect of the random initialization, the clustering for each document set was repeated 50 times, choosing the results with the lowest inertia. To quantify the similarity between clustering results, we used the Normalized Mutual Information (NMI) [23]. For NMI calculation, the average of 5 runs of the aforementioned clustering process was computed. Each clustering was therefore conducted $50\times 5$ times to minimize the impact of random initializations.
@@
Clustering-based utility evaluation for the considered strategies are shown in Figure 6. If we compare the results of this clustering-based utility evaluation with TPS, we obtain a strong Pearson correlation $r=0.864$ ($p$-$value=0.012$). This shows the effectiveness of TPS as an a priori estimator of the analytical utility of the protected documents in downstream tasks. Minor differences between approaches stem from TPS assessing semantic similarity in isolation, without considering the broader linguistic context. A term may be semantically appropriate alone but poorly integrated in its surrounding context. While downstream results suggest minimal impact on the overall semantic content, TPS scores does not directly account for readability and stylistic nuances.
Figure 6: Normalized Mutual Information (NMI) for the clustering results obtained after applying the various replacement strategies (higher is better). The three last results (in green) leverage the Mistral 7B instruct model.
@@
Method-level differences for NMI offer additional insights into how various approaches impact empirical data utility. For instance, INTACT achieves the highest NMI, even higher than the most-specific replacements, while their TPS remains lower. Similarly, entity replacements obtained a higher NMI than least-specific replacements, while the opposite holds for TPS. A common factor in both cases is the use of entity types. Although generated replacements are more similar to the original text spans when examined individually (i.e., higher TPS), some tend to distort the general document content more than entity types (i.e., lower NMI). Such distortion likely arises when a generalization leads to overly broad replacements that disrupt the linguistic register and structure of the document (as exemplified by the three least-specific replacements in Example 1), or even unfaithfully abstract its initial content. The strategy of using entity types when all generated replacements are guessed (which is frequent, as shown in Figure 4) actually resulted in a NMI that exceeded that of the most-specific replacements. It seems that, when the language model fails to generate replacements that sufficiently protect privacy, it also struggles to produce options that effectively preserve utility.
@@
Discrepancy between TPS and NMI is more pronounced with Dou et al. and least-specific replacements, which achieved a TPS nearly identical to INTACT but resulted in a much lower NMI. The replacements obtained with these methods are often lengthy sentences that, while resembling the original span (i.e., high TPS), tend to interfere more with the general content of the text (and thus yielding a low NMI).
@@
Presidio’s synthetic replacements achieved a high NMI, outperforming both entity-based and least-specific replacements, likely due to their better integration within the text. Nonetheless, even when using a more advanced model than the one used to generate our replacements (i.e., GPT-3.5-turbo instead of Mistral-7B), its non-truthful replacements still yielded a lower NMI compared to our most-specific and INTACT’s selected replacements.
@@ 5.6 Privacy evaluation
We evaluated the residual disclosure risk of the strategies by using the TRIR metric introduced in Section 4.2.
To this end, we leveraged the ready-to-use implementation available in the corresponding GitHub repository  [^15] https://github.com/BenetManzanaresSalor/TextRe-Identification with the default hyperparameters, including the use of distilbert-base-uncased [^16] https://huggingface.co/distilbert/distilbert-base-uncased [55] as language model. This model is fine-tuned for the re-identification task, what allows to obtain results competitive with those of pre-trained LLMs [11] without extensive computational resources. For instance, fine-tuning on an Nvidia RTX 4080 took just 36 minutes.
TRIR rests on the definition of a body of background knowledge from which re-identification risk is assessed. To define this background knowledge, we used a corpus of 13,759 rulings published on the ECHR website. Since the annotated court decisions in TAB were drawn from this same source, the full corpus represents a strict superset of the TAB documents. However, assuming that an adversary has access to the entire corpus is unrealistic –such an adversary would already possess the original documents and thus have no incentive to re-identify their sanitized versions. It is also unrealistic to assume the adversary’s knowledge includes only the individuals in the protected dataset, as adversaries typically lack precise knowledge of the dataset’s population. Given these considerations, we defined the background knowledge as a random superset of TAB, comprising twice the number of individuals/documents (i.e., $1,268\times 2=2,536$), and included only the first 10% of non-Procedure text from each ruling.
@@
Figure 7: Text Re-identification Risk (TRIR) obtained with the various replacement strategies (lower is better). Results for our approach (marked in green) leveraged the Mistral 7B instruct model.
As expected, the replacements with the most-specific LLM generalizations yield the highest (i.e. worst) TRIR scores. The difference between this approach and the one yielding the lowest TRIR score (the suppression strategy) is, however, relatively small (2.2 percent points). In other words, the reliance on informative replacements (generalizations) makes it possible to substantially improve utility without substantially impairing privacy with respect to the strictest protection, i.e., suppression. INTACT performed particularly well on this balance, achieving a relatively low TRIR overall (second lowest when considering only generated replacements), while delivering the second-highest TPS and the highest NMI. Compared to both the least-specific and most-specific replacements, INTACT delivered lower TRIR alongside comparable or superior TPS and NMI.
On the other hand, Presidio’s and entity replacements incurred the lowest risks because of the non-truthful replacements of the former (see Section 5.7), and the abstract replacements of the latter tended to confuse the re-identification model. This reduced re-identification accuracy even more than suppression, because the re-identification model relies on concepts found in the background knowledge of the individuals targeted for re-identification. While suppression removes part of that exploitable information, Presidio’s synthetic and entity replacements substitute it by new information that is rarely found in any background knowledge.
The scenario differs for truthful replacements –like INTACT and Dou et al.–, where part of the meaning of the original concept is preserved, making it possible to connect it with the corresponding individual’s background knowledge, which slightly increases TRIR. However, regardless of the actual TRIR, the use of non-truthful replacements, as is the case for e.g. Presidio, may allow an attacker to differentiate between protected and non-protected terms, thereby providing an advantage that can be exploited to conduct case-specific attacks focused on inferring only the protected terms. Furthermore, a document bearing explicit markers of text sanitization (such as suppressions or replacements with generic placeholders) raises awareness of the document’s sensitivity and may thereby attract more attention as a target for attacks [9].
@@ 5.7 Truthfulness and abstraction level
An important characteristic of the INTACT approach is its emphasis on producing truth-preserving replacements whose semantic content subsumes the content of the original text span. Such replacements should therefore correspond to generalizations that encompass and abstract away the original text span. However, the truthfulness and abstraction level of the replacements are nuances that are challenging to evaluate automatically in a reliable manner. Their impact in combination is reflected in the results of our automatic metrics, but these do not provide explicit information about replacement truthfulness and abstraction level separately.
@@
To address this, we conducted a manual evaluation focused on the truthfulness and abstraction level of the selected replacements. This evaluation was performed on a subset of TAB documents, and comprised 600 instances, each consisting of a PII span mention with 70 characters preceding and following it as context. For each document, we randomly sampled a mention for each entity and then selected a random subset of all these sampled mentions, using an equal size of 100 instances per each entity type, excluding the PERSON and CODE categories. The instances were rated by a human evaluator with NLP background, who was also allowed to look up concepts online whenever necessary.
@@
The selected replacements were assessed using two criteria: abstraction level and truthfulness. A replacement was considered more abstract if it constituted a hypernym or provided fewer details than the original PII span. A replacement was truthful if it retained the same overall content of the original span. When approximate equivalence could be assumed in the context of the span (e.g.28 days and 1 month), it was counted as truth-preserving. Partially true spans (e.g., containing two dates) were not considered truthful.
To determine the statistical reliability of this human evaluation, we measured inter-rater agreement using assessments from an additional human rater. We randomly sampled a total of 148 instances for this purpose, equally distributed across the three methods and the six sanitized entity types used. The two raters agreed on 90% of specificity evaluations and 93% of truthfulness evaluations. Cohen’s $\kappa$ [13] which accounts for chance agreement, was 0.76 for specificity and 0.84 for truthfulness, corresponding to substantial and almost perfect agreement, respectively [34]. Differences in annotators’ specificity judgments were observed in cases where proper names in the original span were replaced with common nouns conveying the same meaning (e.g., Widow’s Bereavement Allowance replaced by a specific allowance for widows). Some disagreements regarding truthfulness likely arose from the subjective interpretation of replacements with vague wording (e.g., late evening for 7.30 p.m.) and from numeric information where the original and the replacement spans differred only slightly and were thus considered approximately equivalent (truthful) in some cases (e.g., replacing 10.45 with 10.46). Another potential source of disagreement was polysemy, as in the case of whether a facility for detention is a truthful replacement for Oakington, which refers to both a village and an immigration detention center in the United Kingdom.
@@
Table 2 presents the manual evaluation results for Presidio, Dou et al.’s model [18] and INTACT. According to this manual evaluation, 81.3% of the INTACT replacements were deemed as less specific and 93.2% as truthful. In contrast, Presidio’s manual evaluation reports 36.2% less specific and just 19.7% truthful replacements. Dou et al.’s abstraction model also produced mostly more abstract and mostly truthful replacements (78.3% and 95.2% respectively). However, replacements that were truthful but not more abstract amounted to 20.2% for Dou et al., which is 5% higher than for INTACT, where the iterative inference attacks are meant to prevent the use of too specific replacements. This corroborates the findings from the TRIR-based automatic evaluation results (Figure 7) which show a higher privacy risk posed by Dou et al. than INTACT.
@@
Table 2: Human evaluation of replacements on the truthfulness and abstraction level for three selected sanitization methods: Presidio, Dou et al and INTACT (our approach).
Approach: Not-truthful Truthful
@@
Presidio [39] 61.8% 18.5% 2% 17.7%
Dou et al. [18] 1.5 % 3.3% 20.2% 75%
INTACT 3.5% 3.3% 15.2% 78%
The results from Table 2 show that smaller, open LLMs are well suited to finding generalized replacements in terms of abstraction level and, especially, truthfulness, even without specific fine-tuning for this task. For both INTACT and Dou et al., however, the generation of more abstract replacements was somewhat more challenging for the DEM and the MISC categories, where text spans are often more specialized (typically medical and legal) terms or foreign, possibly culture-specific words. The LLMs showed some tendency to produce definitions, explanations or paraphrases for these terms (e.g., a report on an individual’s health status from a medical institution as a replacement for the Swedish medical term epikris).
In some cases, none of the suggestions were generalizations, e.g., unemployed resulted in the INTACT replacements: Out of work, Jobless, Inactive in the labor force, Economically inactive, Without employment. We also noticed that the truthful and more abstract replacements obtained with Dou et al. were often more generic (e.g., in a specific period for replacing a DATETIME entity) than the ones for INTACT. This type of semantic loss is likely one of the reasons why INTACT outperformed Dou et al. in the clustering-based utility evaluation results in Figure 6.
In comparison, Presidio’s replacements were much less consistent, with a vast majority of them being non-truthful and with arbitrary degrees of abstraction. The lack of truthfulness and consistency is expected to significantly distort the document’s meaning. Moreover, we noticed a substitution bias towards US-related sources, with the nationality of defendants in the TAB corpus changed from Turkish, Polish and British to Brazilian, Canadian and Mexican. Even though these distortions contribute to preventing re-identification (as shown in Section 5.6), they also seriously hamper secondary uses, as it might invalidate conclusions drawn from per-document analysis.